A stock has had returns of 23 percent, 13 percent, 2 percent, -12 percent, -10 percent, and 19 percent over the last six years. What are the arithmetic and geometric returns for the stock? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1

Arithmetic Returns of the stock is simply the mathematical average of the stock's returns over the last 6 years

Therefore, Arithmetic Stock Return = (23 + 13 + 2 + (-12) + (-10) + 19) / 6 = 5.83 %

Geometric Return of Stock = [(1.23) x (1.13) x (1.02) x (1-0.12) x (1-0.1) x (1.19)]^(1/6) - 1 = 0.04948 ~ 4.95 %